The iq of 300 students in a certain school in Quezon province is approximately normally distributed with mean of 100 and standard deviation is 15??


1
Expert's answer
2022-01-06T18:06:34-0500

Sample size (n) = 300

Mean = 100

Standard deviation = 15

The percentage of students have an IQ from 85 to 120 is:

"P(85<X<120)=P(\frac{X-\mu}{\frac{\sigma}{\sqrt{n}}}<Z<\frac{X-\mu}{\frac{\sigma}{\sqrt{n}}})"


"P(85<X<120)=P(\frac{85-100}{\frac{15}{\sqrt{300}}}<Z<\frac{120-100}{\frac{15}{\sqrt{300}}})"


"P(85<X<120)=P(-17.32<Z<23.09)"


"(85<X<120)=P(Z<23.09)-(Z<-17.32)"


Find probability in respect to each Z score using standard distribution table.


"(85<X<120)=1.0-0.00=1.0\approx100\%"


This means approximately 100% students have an IQ from 85 to 120.